Circular No. 8285

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2002 (121) 1 J. Lecacheux, Observatoire de Paris, reports that recent observations of

the satellite of (121) Hermione (cf. IAUC 7980) — obtained during Jan. 4– 6 with the Very Large Telescope and Keck Observatory adaptative-optics system by F. Marchis, I. De Pater, D. Hestroffer, P. Descamps, and J. Berthier — have been used to compute the probable ground track of an occultation predicted to be visible for southern-European observers for up to ∼ 3 s around Feb. 16.938 UT. The parent object (121) will occult the same ninth-magnitude star, BD +29^{o}1354 (α = 6^{h}49^{m}15^{s}, δ = +29^{o}06′.2, equinox 2000.0), for ∼ 35 s along a parallel path of visibility located ∼ 1^{o}.55 south and 64 ± 5 s earlier [see the prediction by Dunham 2004, Sky Tel. 107(3), 102] with respect to the occultation track for the satellite. The retrograde orbit of S/2002 (121) 1 about the primary (i ∼ 150^{o}, P = 1.62722 ± 0.00007 days) seems to be precessing by ∼ +40^{o}/yr around a pole located within the approximate box α = [10^{o}, 20^{o}], δ = [−10^{o}, 0^{o}]. On Feb. 16.938, the satellite is predicted to be at Δα = +0^{s}.0137 ± 0^{s}.0010, Δδ = +0′′.081 ± 0′′.018 with respect to the parent body. The 18-km-wide shadow of S/2002 (121) 1 is predicted to pass across northern Greece, Albania, central Italy, and southern France (at long. 34^{o}.793 E, lat. 36^{o}.045 N on Feb. 16.9347 and at long. 4^{o}.722 W, lat. 45^{o}.230 N on Feb. 16.9410; estimated absolute uncertainties are ± 75 km for lateral shift and ± 12 s of time). Updates and details can be obtained at website URL http://www.euraster.net.

SUPERNOVAE 2004T, 2004U, AND 2004V Further to IAUC 8281, J. Graham, B. Swift, and W. Li report the

LOSS discovery of three apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004T Feb. 9.4 10^{h}56^{m}10.^{s}27 +47^{o}23′43′′.5 17.5 6′′.1 E, 10′′.7 S
2004U Feb. 10.5 11 07 18.83 +28 22 23.6 17.7 1′′.7 E, 0′′.9 S
2004V Feb. 10.6 14 50 55.33 +17 13 12.6 18.1 1′′.2 W, 7′′.5 N

The offset for SN 2004U is with respect to the center of a galaxy that itself is 6′′.7 east and 44′′.5 north of NGC 3527. Additional unfiltered KAIT magnitudes: SN 2004T in UGC 6038, Jan. 23.4, [19.0; Feb. 10.4, 17.5. SN 2004U, 2003 May 13.2, [19.0; 2004 Jan. 21.5, 17.9. SN 2004V, Jan. 22.6, [19.0, Feb. 11.6, 18.0.

2004 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green